# Sequential Monte-Carlo testing by betting

### v4

@@ 5.2 Uniform prior: closed-form wealth is monotone in losses
The wealth of the binomial mixture strategy with density $u_{c}$ after $B$ permutations and $k$ losses is given by $\bar{W}_{B}^{u_{c}}(k)=(1-\mathrm{Bin}(k;B+1,c))/c$, where $\mathrm{Bin}(\cdot;B+1,c)$ denotes the cumulative distribution function of a binomial distribution with $B+1$ trials of probability $c$.
@@
for all $B\in\mathbb{N}$ and $k\in\{0,1,\ldots,B\}$ by (backward) induction in $k$.
@@
=$\stackrel{{\scriptstyle\text{IH}}}{{=}}{B+1\choose k}c^{k}(1-c)^{B-k+1}+\sum_{i=k+1}^{B+1}c^{i}(1-c)^{B-i+1}{B+1\choose i}$
$=\sum_{i=k}^{B+1}c^{i}(1-c)^{B-i+1}{B+1\choose i}.$
@@
5.3 Uniform prior: zero resampling risk
The wealth of the above binomial mixture strategy has many desirable properties. For example, in contrast to the simple binomial strategy, the wealth is decreasing for an increasing number of losses. This is shared with the classical permutation p-value and thus seems intuitive. In the following, we show that this binomial mixture strategy also shares some asymptotic behavior with the permutation p-value, which has implications for its resampling risk.
@@ 5.3 Uniform prior: zero resampling risk
$\bar{W}_{B}^{u_{c}}|\{\mathfrak{p}^{\mathrm{lim}}=p^{\mathrm{lim}}\}\stackrel{{\scriptstyle a.s.}}{{\to}}\begin{cases}1/c,&\text{ if }p^{\mathrm{lim}}\in[0,c)\\ 0,&\text{ if }p^{\mathrm{lim}}\in(c,1]\end{cases}$
for $B\to\infty$.
@@
Let $p^{\mathrm{lim}}\in[0,c)$. By the De Moivre-Laplace Theorem the binomial distribution approaches a normal distribution, meaning for every $\epsilon>0$ there exists $B_{1}\in\mathbb{N}$ such that $\mathrm{Bin}(k;B+1,c)\leq\Phi((k-Bc)/(\sqrt{Bc(1-c)}))+\epsilon$ for all $B\geq B_{1}$ and $k\in\{0,\ldots,B\}$, where $\Phi$ is the cumulative distribution function of a standard normal distribution. In addition, since $\nicefrac{{K_{B}}}{{B}}|\{\mathfrak{p}^{\mathrm{lim}}=p^{\mathrm{lim}}\}\stackrel{{\scriptstyle a.s.}}{{\to}}p^{\mathrm{lim}}$ for $B\to\infty$ and $p^{\mathrm{lim}}<c$, there exists a $B_{2}\in\mathbb{N}$ a.s. such that $K_{B}|\{\mathfrak{p}^{\mathrm{lim}}=p^{\mathrm{lim}}\}\leq B(p^{\mathrm{lim}}+\delta)$ for $B\geq B_{2}$ and some $\delta>0$ with $\delta+p^{\mathrm{lim}}<c$. Hence,
$\mathrm{Bin}(K_{B};B+1,c)|\{\mathfrak{p}^{\mathrm{lim}}=p^{\mathrm{lim}}\}\leq\Phi\left(\frac{\sqrt{B}(p^{\mathrm{lim}}+\delta-c)}{\sqrt{c(1-c)}}\right)+\epsilon$
for all $B\geq\max(B_{1},B_{2})$, which goes to $\epsilon$ for $B\to\infty$. Analogously, it follows that $\mathrm{Bin}(K_{B};B+1,c)|\{\mathfrak{p}^{\mathrm{lim}}=p^{\mathrm{lim}}\}\stackrel{{\scriptstyle a.s.}}{{\to}}1$ if $p^{\mathrm{lim}}\in(c,1]$. ∎
Theorem 5.3 particularly shows that for $c=\alpha$, the wealth $\bar{W}_{B}^{u_{\alpha}}|\{\mathfrak{p}^{\mathrm{lim}}=p^{\mathrm{lim}}\}\stackrel{{\scriptstyle\text{a.s.}}}{{\to}}1/\alpha$ if $p^{\mathrm{lim}}<\alpha$ and $\bar{W}_{B}^{u_{\alpha}}|\{\mathfrak{p}^{\mathrm{lim}}=p^{\mathrm{lim}}\}\stackrel{{\scriptstyle\text{a.s.}}}{{\to}}0$ if $p^{\mathrm{lim}}>\alpha$ for $B\to\infty$. Hence, if the event that the limiting permutation p-value equals $\alpha$ is a null set, meaning $\mathbb{P}(\mathfrak{p}^{\mathrm{lim}}=\alpha)=0$, then $\bar{W}_{B}^{u_{\alpha}}$ converges to the all-or-nothing bet [26] on $\mathfrak{p}^{\mathrm{lim}}$, $1\{\mathfrak{p}^{\mathrm{lim}}\leq\alpha\}/\alpha$, almost surely. However, note that with the choice $c=\alpha$, Proposition 5.2 implies that $\bar{W}_{B}^{u_{\alpha}}<1/\alpha$ for all $B\in\mathbb{N}$ (meaning that if it converges to $1/\alpha$, this convergence occurs from below), and so we could never stop after a finite number of permutations. For this reason, when aiming for a level-$\alpha$ test in practice, one should choose $c$ slightly smaller than $\alpha$, which ensures a wealth of $\bar{W}_{B}^{u_{c}}>1/\alpha$ after a finite number of permutations if $p^{\mathrm{lim}}<c$. This is captured in the following corollary, which should be compared to Corollary 4.5.
@@
Let $c<\alpha$ and $\tau:=\inf\{t\geq 1:\bar{W}_{t}^{u_{c}}\geq 1/\alpha\}$. If $\mathfrak{p}^{\mathrm{lim}}\in[0,c)$, then $\tau$ is almost surely finite and the test $\phi=1\{\bar{W}_{\tau}^{u_{c}}\geq 1/\alpha\}$ has zero resampling risk ($RR_{p^{\mathrm{lim}}}(\phi)=0$).
The parameter $c<\alpha$ can also be interpreted as the aggressiveness with which we bet on the alternative. If we choose $c$ very close to $\alpha$, we have the guarantee to always reject if $p^{\mathrm{lim}}\leq\alpha-\epsilon$ for an arbitrary small $\epsilon>0$. However, if the difference between $\alpha$ and $c$ is larger, we might be able to reject much earlier, which can be used to save computational cost. Thus, the binomial mixture strategy with density $u_{c}$ offers a simple trade-off between power and computation time.
The closed form of the wealth in Proposition 5.2 can easily be used to calculate or bound the required number of permutations for concrete parameter constellations. For example, if we observe no losses, we can reject the hypothesis as soon as $(1-(1-c)^{B+1})/c>1/\alpha$. Hence, if $\alpha=0.05$ and $c=0.04$, we can reject the null hypothesis after $39$ permutations and if $c=0.049$, we can reject after $77$ permutations. As another example, one can calculate that $\bar{W}_{200}^{u_{0.04}}(5)>1/\alpha$. Thus, if we choose $c=0.04$, we can reject the null hypothesis at level $\alpha=0.05$ after $\leq 200$ permutations if we observe $\leq 5$ losses during these permutations.
@@ 5.4 Stochastic rounding of promising wealth
Suppose we stopped the testing process at some time $\tau$ where the current wealth looks promising ($W_{\tau}$ close to $1/\alpha$) but did not lead to a rejection yet ($\max_{s=1,\ldots,\tau}W_{s}<1/\alpha$). For example, this might be when a prespecified maximum number of permutations is hit or the wealth changed extremely slowly during the last permutations. As described in Section 1.1, we could continue sampling or interpret the wealth on its own, since a large wealth is still good evidence against the null hypothesis. If both these options fall out, as drawing more permutations is not possible or wanted and we are only interested in rejections, we could also use a technique called stochastic rounding to achieve (possibly) a rejection [37].
This is based on a randomized improvement of Ville’s inequality by Ramdas and Manole [19]. They showed that for any stopping time $\tau$ with respect to the filtration $(\mathcal{R}_{t})_{t\geq 1}$, it holds that
@@
where $\alpha\in(0,1)$ and $U\sim U[0,1]$ is independent of the filtration $(\mathcal{R}_{t})_{t\geq 1}$. Thus, if we stopped at time $\tau$ before reaching the level $1/\alpha$, we can draw a sample $u$ from a uniformly distributed random variable $U\sim U[0,1]$ and reject the null hypothesis if $W_{\tau}\geq u/\alpha$. However, in order to maintain validity, it is important that $u$ is sampled after stopping the process, such that $\tau$ and $W_{\tau}$ are independent of $U$, and that it is only sampled once (and not multiple times to pick the best one). Therefore, this technique should only be used in an automated code which prevents this randomized improvement from being exploited for cheating. The term stochastic rounding stems from this test being equivalent to rejecting if $W_{\tau}^{\sim}\geq 1/\alpha$, where $W_{\tau}^{\sim}$ is an e-value defined by $W_{\tau}^{\sim}=W_{\tau}$ if $W_{\tau}\geq 1/\alpha$; and $W_{\tau}^{\sim}=1/\alpha$ with probability $W_{\tau}\alpha$ and $W_{\tau}^{\sim}=0$ with probability $(1-W_{\tau})\alpha$ if $W_{\tau}<1/\alpha$.
For example, stochastic rounding could be used to set the resampling risk for all $p^{\mathrm{lim}}<\alpha$ to an arbitrary small $\epsilon\in(0,1)$ based on the binomial mixture strategy with prior $u^{\alpha}$, which follows immediately by Theorem 5.3.
@@
Let $\epsilon\in(0,1)$ and $\tau=\inf\{t\geq 1:\bar{W}_{B}^{u_{\alpha}}\geq(1-\epsilon)/\alpha\}$. If $\mathfrak{p}^{\mathrm{lim}}\in[0,\alpha)$, then $\tau$ is almost surely finite and the randomized test $\phi=1\{\bar{W}_{\tau}^{u_{\alpha}}\geq(1-\epsilon)/\alpha,U\leq(1-\epsilon)\}$ has resampling risk $\epsilon$ ($\mathrm{RR}_{p^{\mathrm{lim}}}(\phi)=\epsilon$).
@@ 6 Every permutation test can be recovered by betting
A common concern regarding sequential testing and the use of e-values is that power is lost compared to classical p-value methods like the permutation p-value $\mathfrak{p}^{\mathrm{perm}}$. In this section, we show that every permutation test that is a function of the rank of $Y_{0}$ after a fixed number of permutations can be obtained by our strategy. In particular, we demonstrate that the permutation p-value and the Besag-Clifford method arise naturally as concrete instances of our general algorithm (Algorithm 2) by prespecifying the number of permutations or the number of losses, respectively.
Consider an e-value $E_{t}(\mathrm{rank}(Y_{0};Y_{0}^{t}))$ that is defined by a non-random function $E_{t}$ of the rank of the observed test statistics $\mathrm{rank}(Y_{0};Y_{0}^{t})$ for some prespecified $t\geq 1$. Similarly to $B_{t}$, $E_{t}$ can be interpreted as a $t+1$-dimensional betting vector where $E_{t}(r)$ is our realized wealth if $\mathrm{rank}(Y_{0};Y_{0}^{t})=r$. Instead of betting sequentially at each step, the e-value $E_{t}(\mathrm{rank}(Y_{0};Y_{0}^{t}))$ is just obtained by betting once on the rank of the observed test statistic $Y_{0}$. Analogously to $B_{t}$, in order to define an (admissible) e-value the betting vector $E_{t}$ needs to satisfy
$\sum_{r=1}^{t+1}E_{t}(r)=t+1.$
We first show that every $E_{t}$ can also be obtained by our sequential betting strategy — this implies that we do not lose something due to the sequential betting. For this, note that after betting with $E_{t-1}$ on the rank of $Y_{0}$ among $Y_{0},\ldots,Y_{t-1}$, we could also draw a further permuted test statistic $Y_{t}$, bet on its rank among $Y_{0},\ldots,Y_{t}$ with $B_{t}$ and define a new e-value as $E_{t-1}(\mathrm{rank}(Y_{0};Y_{0}^{t-1}))\cdot B_{t}(R_{t})$. We claim that every e-value $E_{t}(\mathrm{rank}(Y_{0};Y_{0}^{t}))$ can be obtained by such a product. Given an arbitrary betting vector $E_{t}$, this problem can be transferred into the following system of equations:

### v5

@@ 5.2 Uniform working prior: closed-form wealth
The wealth of the binomial mixture strategy with density $u_{c}$ after $T$ permutations and $\ell$ losses is given by $\bar{W}_{T}^{u_{c}}(\ell)=(1-\mathrm{Bin}(\ell;T+1,c))/c$, where $\mathrm{Bin}(\cdot;T+1,c)$ denotes the cumulative distribution function of a binomial distribution with $T+1$ trials of probability $c$.
@@
for all $T\in\mathbb{N}$ and $\ell\in\{0,1,\ldots,T\}$ by (backward) induction in $\ell$.
@@
=$\stackrel{{\scriptstyle\text{IH}}}{{=}}{T+1\choose\ell}c^{\ell}(1-c)^{T-\ell+1}+\sum_{i=\ell+1}^{T+1}c^{i}(1-c)^{T-i+1}{T+1\choose i}$
$=\sum_{i=\ell}^{T+1}c^{i}(1-c)^{T-i+1}{T+1\choose i}.$
@@
With this, we can apply the binomial mixture strategy with working prior $u_{c}$ by just using the closed-form wealth at each step (see Algorithm 3). In the next subsection, we argue why we usually choose $c$ smaller than, but close to $\alpha$. One might think that a beta distribution with decreasing density is a better working prior for $Q_{\mathrm{lim}}$ than $u_{c}$ and therefore should yield a higher power. In Appendix E, we demonstrate via simulations that this is not the case.
Algorithm 3 The binomial mixture strategy with working prior $u_{c}$, along with stopping for futility
Input: Significance level $\alpha\in(0,1)$, parameter $c<\alpha$ and sequence of test statistics $Y_{0},Y_{1},Y_{2},\ldots$
      Output: Stopping time $\tau$ and wealth $\bar{W}_{1}^{u_{c}},\ldots,\bar{W}_{\tau}^{u_{c}}$.
$\bar{W}_{0}^{u_{c}}=1$
2:$L_{0}=0$
3:for $t=1,2,...$ do
4:     if $Y_{t}\geq Y_{0}$ then
5:         $L_{t}=L_{t-1}+1$
6:     else
7:         $L_{t}=L_{t-1}$
8:     end if
9:     $\bar{W}_{t}^{u_{c}}=(1-\mathrm{Bin}(L_{t};t+1,c))/c$
10:     if $\bar{W}_{t}^{u_{c}}<\alpha\textbf{ or }\bar{W}_{t}^{u_{c}}\geq 1/\alpha$ then
11:         $\tau=t$
12:         return $\tau,\bar{W}_{1}^{u_{c}},\ldots,\bar{W}_{\tau}^{u_{c}}$
13:     end if
14:end for
5.3 Uniform working prior: zero resampling risk
The wealth of the above binomial mixture strategy has many desirable properties. For example, in contrast to the simple binomial strategy, the wealth is decreasing for an increasing number of losses. In Section 3, we noted that this is actually the case for all working priors yielding a distribution stochastically smaller than a uniform on $[0,1]$. This monotonicity in the number of losses is shared with the classical permutation p-value and thus seems intuitive. In the following, we show that the binomial mixture strategy with working prior $u_{c}$ also shares some asymptotic behavior with the permutation p-value, which has implications for its resampling risk.
@@ 5.3 Uniform working prior: zero resampling risk
$\bar{W}_{T}^{u_{c}}|\{Q_{\mathrm{lim}}=q_{\mathrm{lim}}\}\stackrel{{\scriptstyle a.s.}}{{\to}}\begin{cases}1/c,&\text{ if }q_{\mathrm{lim}}\in[0,c)\\ 0,&\text{ if }q_{\mathrm{lim}}\in(c,1]\end{cases}$
for $T\to\infty$.
@@
Let $q_{\mathrm{lim}}\in[0,c)$. By the De Moivre-Laplace Theorem the binomial distribution approaches a normal distribution, meaning for every $\epsilon>0$ there exists $T_{1}\in\mathbb{N}$ such that $\mathrm{Bin}(\ell;T+1,c)\leq\Phi((\ell-Tc)/(\sqrt{Tc(1-c)}))+\epsilon$ for all $T\geq T_{1}$ and $\ell\in\{0,\ldots,T\}$, where $\Phi$ is the cumulative distribution function of a standard normal distribution. In addition, since $\nicefrac{{L_{T}}}{{T}}|\{Q_{\mathrm{lim}}=q_{\mathrm{lim}}\}\stackrel{{\scriptstyle a.s.}}{{\to}}q_{\mathrm{lim}}$ for $T\to\infty$ and $q_{\mathrm{lim}}<c$, there exists a $T_{2}\in\mathbb{N}$ such that $L_{T}|\{Q_{\mathrm{lim}}=q_{\mathrm{lim}}\}\leq T(q_{\mathrm{lim}}+\delta)$ a.s. for $T\geq T_{2}$ and some $\delta>0$ with $\delta+q_{\mathrm{lim}}<c$. Hence,
$\mathrm{Bin}(L_{T};T+1,c)|\{Q_{\mathrm{lim}}=q_{\mathrm{lim}}\}\leq\Phi\left(\frac{\sqrt{T}(q_{\mathrm{lim}}+\delta-c)}{\sqrt{c(1-c)}}\right)+\epsilon$
for all $T\geq\max(T_{1},T_{2})$, which goes to $\epsilon$ for $T\to\infty$. Analogously, it follows that $\mathrm{Bin}(L_{T};T+1,c)|\{Q_{\mathrm{lim}}=q_{\mathrm{lim}}\}\stackrel{{\scriptstyle a.s.}}{{\to}}1$ if $q_{\mathrm{lim}}\in(c,1]$. ∎
Theorem 5.3 particularly shows that for $c=\alpha$, the wealth $\bar{W}_{T}^{u_{\alpha}}|\{Q_{\mathrm{lim}}=q_{\mathrm{lim}}\}\stackrel{{\scriptstyle\text{a.s.}}}{{\to}}1/\alpha$ if $q_{\mathrm{lim}}<\alpha$ and $\bar{W}_{T}^{u_{\alpha}}|\{Q_{\mathrm{lim}}=q_{\mathrm{lim}}\}\stackrel{{\scriptstyle\text{a.s.}}}{{\to}}0$ if $q_{\mathrm{lim}}>\alpha$ for $T\to\infty$. Hence, if the event that the limiting permutation p-value equals $\alpha$ is a null set, meaning $\mathbb{P}(Q_{\mathrm{lim}}=\alpha)=0$, then $\bar{W}_{T}^{u_{\alpha}}$ converges to the all-or-nothing bet [31] on $Q_{\mathrm{lim}}$, $\mathbbm{1}\{Q_{\mathrm{lim}}\leq\alpha\}/\alpha$, almost surely. However, note that with the choice $c=\alpha$, Proposition 5.2 implies that $\bar{W}_{T}^{u_{\alpha}}<1/\alpha$ for all $T\in\mathbb{N}$ (meaning that if it converges to $1/\alpha$, this convergence occurs from below), and so we could never stop after a finite number of permutations. For this reason, when aiming for a level-$\alpha$ test in practice, one should choose $c$ slightly smaller than $\alpha$, which ensures a wealth of $\bar{W}_{T}^{u_{c}}>1/\alpha$ after a finite number of permutations if $q_{\mathrm{lim}}<c$. This is captured in the following corollary, which should be compared to Corollary 4.5.
@@
Let $c<\alpha$ and $\tau:=\inf\{t\geq 1:\bar{W}_{t}^{u_{c}}\geq 1/\alpha\}$. If $Q_{\mathrm{lim}}\in[0,c)$, then $\tau$ is almost surely finite and the test $\phi=\mathbbm{1}\{\bar{W}_{\tau}^{u_{c}}\geq 1/\alpha\}$ has zero resampling risk ($RR_{q_{\mathrm{lim}}}(\phi)=0$).
The parameter $c<\alpha$ can also be interpreted as the aggressiveness with which we bet on the alternative. If we choose $c$ very close to $\alpha$, we have the guarantee to always reject if $q_{\mathrm{lim}}\leq\alpha-\epsilon$ for an arbitrary small $\epsilon>0$. However, if the difference between $\alpha$ and $c$ is larger, we might be able to reject much earlier, which can be used to save computational cost. Thus, the binomial mixture strategy with density $u_{c}$ offers a simple trade-off between power and computation time.
The closed form of the wealth in Proposition 5.2 can easily be used to calculate or bound the required number of permutations for concrete parameter constellations. For example, if we observe no losses, we can reject the hypothesis as soon as $(1-(1-c)^{T+1})/c>1/\alpha$. Hence, if $\alpha=0.05$ and $c=0.04$, we can reject the null hypothesis after $39$ permutations and if $c=0.049$, we can reject after $77$ permutations. As another example, one can calculate that $\bar{W}_{200}^{u_{0.04}}(5)>1/\alpha$. Thus, if we choose $c=0.04$, we can reject the null hypothesis at level $\alpha=0.05$ after $\leq 200$ permutations if we observe $\leq 5$ losses during these permutations.
To summarize, the binomial mixture strategy with working prior $u_{c}$ allows to specify a parameter $c<\alpha$ arbitrarily close to $\alpha$ such that $H_{0}$ will be rejected almost surely at some finite time if $Q_{\mathrm{lim}}<c$. Note that this guarantee is not possible with the permutation p-value $Q^{\mathrm{perm}}$ nor Besag-Clifford p-value $Q^{\mathrm{BC}}$. No matter how large we choose the maximum number of permutations $T$ and the maximum number of losses $h$, there will always be a positive probability for accepting $H_{0}$. Furthermore, if there is sufficient evidence against $H_{0}$ early in the sampling process, the binomial mixture allows to stop early for rejection, while the permutation p-value and Besag-Clifford method never stop early for rejection.
@@ 5.4 Stochastic rounding of promising wealth
Suppose we stopped the testing process at some time $\tau$ where the current wealth looks promising ($W_{\tau}$ close to $1/\alpha$) but did not lead to a rejection yet ($\max_{s=1,\ldots,\tau}W_{s}<1/\alpha$). For example, this might be when a prespecified maximum number of permutations is hit or the wealth changed extremely slowly during the last permutations. As described in Section 1.1, we could continue sampling or interpret the wealth on its own, since a large wealth is still good evidence against the null hypothesis. If both these options fall out, as drawing more permutations is not possible or wanted and we are only interested in rejections, we could also use a technique called stochastic rounding to possibly achieve a rejection [44].
This is based on a randomized improvement of Ville’s inequality by Ramdas and Manole [24]. They showed that for any stopping time $\tau$ with respect to the filtration $(\mathcal{I}_{t})_{t\geq 1}$, it holds that
@@
where $\alpha\in(0,1)$ and $U\sim U[0,1]$ is independent of the filtration $(\mathcal{I}_{t})_{t\geq 1}$. Thus, if we stopped at time $\tau$ before reaching the level $1/\alpha$, we can draw a sample $u$ from a uniformly distributed random variable $U\sim U[0,1]$ and reject the null hypothesis if $W_{\tau}\geq u/\alpha$. However, in order to maintain validity, it is important that $u$ is sampled after stopping the process, such that $\tau$ and $W_{\tau}$ are independent of $U$, and that it is only sampled once (and not multiple times to pick the best one). Therefore, this technique should only be used in an automated code which prevents this randomized improvement from being exploited for cheating. The term stochastic rounding stems from this test being equivalent to rejecting if $W_{\tau}^{\sim}\geq 1/\alpha$, where $W_{\tau}^{\sim}$ is an e-value defined by $W_{\tau}^{\sim}=W_{\tau}$ if $W_{\tau}\geq 1/\alpha$; and $W_{\tau}^{\sim}=1/\alpha$ with probability $W_{\tau}\alpha$ and $W_{\tau}^{\sim}=0$ with probability $(1-W_{\tau})\alpha$ if $W_{\tau}<1/\alpha$.
For example, stochastic rounding could be used to set the resampling risk for all $q_{\mathrm{lim}}<\alpha$ to an arbitrary small $\epsilon\in(0,1)$ based on the binomial mixture strategy with working prior $u^{\alpha}$, which follows immediately by Theorem 5.3.
@@
Let $\epsilon\in(0,1)$ and $\tau=\inf\{t\geq 1:\bar{W}_{t}^{u_{\alpha}}\geq(1-\epsilon)/\alpha\}$. If $Q_{\mathrm{lim}}\in[0,\alpha)$, then $\tau$ is almost surely finite and the randomized test $\phi=\mathbbm{1}\{\bar{W}_{\tau}^{u_{\alpha}}\geq(1-\epsilon)/\alpha,U\leq(1-\epsilon)\}$ has resampling risk $\epsilon$ ($\mathrm{RR}_{q_{\mathrm{lim}}}(\phi)=\epsilon$).
@@ 6 Every permutation test can be recovered by betting
A common concern regarding sequential testing and the use of e-values is that power is lost compared to classical p-value methods like the permutation p-value $Q^{\mathrm{perm}}$. In this section, we show that every permutation test that is a function of the number of losses $L_{t}$ after a fixed number of permutations $t$ can be obtained by our strategy. In particular, we demonstrate that the permutation p-value and the Besag-Clifford method arise naturally as concrete instances of our general algorithm (Algorithm 1) by prespecifying the number of permutations or the number of losses, respectively.
6.1 Recursive betting strategy for the construction of any permutation e-value
Consider an e-value $E_{t}(L_{t})$ that is defined by a non-random function $E_{t}$ of the number of losses $L_{t}$ for some prespecified $t\geq 1$. Similarly to $B_{t}$, $E_{t}$ can be interpreted as a betting function where $E_{t}(\ell)$ is our realized wealth if $L_{t}=\ell$. Instead of betting sequentially at each step, the e-value $E_{t}(L_{t})$ is just obtained by betting once on the number of losses after $t$ steps. Under $H_{0}$, the number of losses $L_{t}$ are distributed uniformly on $0,1,\dots,t$. Hence, in order to define an (admissible) e-value the betting function $E_{t}$ needs to satisfy
$\sum_{\ell=0}^{t}E_{t}(\ell)=t+1.$ (12)
We first show that every $E_{t}$ can also be obtained by our sequential betting strategy — this implies that we do not lose something due to the sequential betting. For this, note that after betting with $E_{t-1}$ on $L_{t-1}$, we could also draw a further permuted test statistic $Y_{t}$, bet on the indicator $I_{t}=\mathbbm{1}\{Y_{t}\geq Y_{0}\}$ with $B_{t}$ and define a new e-value as $E_{t-1}(L_{t-1})\cdot B_{t}(I_{t})$.
We claim that every e-value $E_{t}(L_{t})$ can be obtained by such a product. Indeed, given an arbitrary betting vector $E_{t}$, and letting $B_{t|\ell}(1)$ (or $B_{t|\ell}(0)$) denote our bet on a loss (or win) at step $t$ if $L_{t-1}=\ell$, consider the following system of equations: